OpenStudy (anonymous):

Convert to trigonometric form. 2 + 2i sqrt 3

OpenStudy (anonymous):

\[2 + 2i \sqrt{3}\]

OpenStudy (noorik.):

indeed, it'll help him

OpenStudy (anonymous):

The answer I got is (4)e^i(2/3)π

OpenStudy (anonymous):

Is that correct?

OpenStudy (anonymous):

\[\Large 2+2 i \sqrt{3}=4 \left(\cos \left(\frac{\pi }{3}\right)+i \sin \left(\frac{\pi }{3}\right)\right)=4 e^{\frac{i \pi }{3}} \]

OpenStudy (anonymous):

Oh man, I was way off.

OpenStudy (anonymous):

\[ 2+2 i \sqrt{3}=4 \left(\frac{1}{2}+\frac{i \sqrt{3}}{2}\right)=4 \left(\cos \left(\frac{\pi }{3}\right)+i \sin \left(\frac{\pi }{3}\right)\right)=4 e^{\frac{i \pi }{3}} \]

OpenStudy (anonymous):

Do you understand it?

OpenStudy (anonymous):

Sort of.